OpenStudy (anonymous):

MAX AND MIN: f(x) = x^3 - 3x + 1 i=[-3/2 , 3]

OpenStudy (cwrw238):

do u know calculus?

OpenStudy (cwrw238):

assuming u do f'(x) =3 x^2 - 3 = 0 for critical points x^2 = 1 x = 1 or -1 we now need to know which is max and which is a minm f"(x) = 6x when x = 1 f"(x) is +ve so this is minm x = -1 is a maxm f(x) maxm when f(x) = 3 minm at f(x) = -1

OpenStudy (cwrw238):

these are local maxm / minm values

OpenStudy (anonymous):

that i represents an interval you're looking at for the function?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i got max = 19

OpenStudy (anonymous):

whenever you look for max/min on a closed interval you must evaluate at the enpoints in addition to the results posted by @cwrw238 . did you get that 19 from the critical points?

OpenStudy (anonymous):

test pts

OpenStudy (anonymous):

how did you get the test points?

OpenStudy (anonymous):

wait, yes i got 19 from my critical its

OpenStudy (anonymous):

oh... ok... did you evaluate at x=-3/2 and x=3 also? because you need to compare that with the 19.

OpenStudy (anonymous):

evaluate the function, not the derivative.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

because this is a closed interval, you should have a minimum also.

OpenStudy (anonymous):

min = -1 at x=1

OpenStudy (anonymous):

i guess you're done... good job. :)

OpenStudy (anonymous):

another q- The sum of 2 nonnegative numbers is 36. Find these numbers if the 1st plus the square of the 2nd is a maximum

OpenStudy (anonymous):

please help me.. my teacher is a awful explainer

OpenStudy (anonymous):

HELP